OpenStudy (anonymous):

The range of f(x) = -1/4 sin 4x is (a) -1/4 ≤ y ≤ 0 (b) -1/4 ≤ y ≤ 1/4 (c) 0 ≤ y ≤1/4 (d) -1 ≤ y ≤ 1 (e) -4 ≤ y ≤ 4

OpenStudy (anonymous):

the range of sine is \(-1\leq y\leq 1\) so what must the range of \(-\frac{1}{4}\sin(x)\) be ?

OpenStudy (anonymous):

is it -1/4 ≤ y ≤ 1/4 ? i'm not sure

OpenStudy (helder_edwin):

yes. u r right

OpenStudy (anonymous):

but in the question there's a 4 in front of the x

OpenStudy (anonymous):

sin4x

OpenStudy (helder_edwin):

check this out http://www.wolframalpha.com/input/?i=sin+x%2C+sin+4x

OpenStudy (helder_edwin):

do u see the difference between sin x and sin 4x

OpenStudy (anonymous):

yeah sin4x has a higher frequency ?

OpenStudy (helder_edwin):

exactly. that's the only difference both oscilate between -1 and 1, though.

OpenStudy (anonymous):

so would the answer be (b) or (d) ?

OpenStudy (helder_edwin):

(b)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (helder_edwin):

u r w